what is the answer to the equation: 13x squared +1=6x
Answer by stanbon(75887)

13x squared +1=6x
----
Rearrange:
13x^2 - 6x + 1 = 0
Use the quadratic formula:
x = [6 +- sqrt(36-4*13*1)]/26
---
x = [6 +- sqrt(-16)]/26
---
x = [6 +- 4i]/26
---
x = (3/13) + (2/13)i or x = (3/13) - (2/13)i
--------------------
Cheers,
Stan H.